OpenStudy (anonymous):

What is the exact diameter of an area rug with a circumference of feet? feet feet feet feet

OpenStudy (anonymous):

Assuming a circular rug, say it is a circumference of x feet, then x = 2*pi* r = pi * diameter so diameter = x/pi